Sunset rents an SUV at $21.95 plus $0.23 per mile. Sunrise rents the same vehicle for $24.95 plus $0.19 per mile. For what mileage is the cost the same?
Hint: Using "m" for miles, write one equation for each rental company. Then solve the system by one of the methods learned in assignment 5.6 or 5.7.

Answer by ankor@dixie-net.com(12701)

Sunset rents an SUV at $21.95 plus $0.23 per mile. Sunrise rents the same vehicle for $24.95 plus $0.19 per mile. For what mileage is the cost the same?
Hint: Using "m" for miles, write one equation for each rental company.
:
Sunset cost expression: .23m + 21.95
Sunrise cost expression: .19m + 24.95
:
Sunset cost = Sunrise cost, at what mileage (m)
.23m + 21.95 = .19m + 24.95
:
Perform the necessary operations to get the variable (m) on the left by itself:
.23m - .19m = 24.95 - 21.95
.04m = 3.00
m = 3/.04
m = 75 miles when the two rental plans are equal
:
:
Check,substitute 75 for x:
Sunset: .23(75) + 21.95 = 39.20
Sunrise: .19(75) + 24.95 = 39.20